OpenStudy (vayacondios1):

In ΔABC, if the length of side b is 3 centimeters and the measures of angle B and angle C are 45° and 60°, respectively, what is the length of side c to two decimal places? 2.45 3.67 4.00 5.45 will fan helpers

OpenStudy (vayacondios1):

no idea how this works

OpenStudy (peachpi):

you can use the law of sines \[\frac{sinA}{a}=\frac{ \sin B }{ b }=\frac{ sinC }{ c }\]

jhonyy9 (jhonyy9):



OpenStudy (vayacondios1):

is it 75

OpenStudy (vayacondios1):

degrees?

OpenStudy (peachpi):

A is 75°, but you don't really need it. Just use the B and C parts of the equation. You know that b = 3 cm, B = 45°, and C = 60°. The only thing missing is c.

OpenStudy (peachpi):



OpenStudy (vayacondios1):

how would you find that?

OpenStudy (peachpi):

plug in the numbers \[\frac{ sin B }{ b }=\frac{ sin C }{ c }\]

OpenStudy (vayacondios1):

ill try

OpenStudy (vayacondios1):

i cant figure it out sorry

OpenStudy (peachpi):

let me see what you have

OpenStudy (vayacondios1):

i couldnt find the sin of b or sin c

OpenStudy (peachpi):

uppercase letters are angles, lowercase are sides. Given in the question: b = 3 cm, B = 45°, and C = 60°. plug into a calculator to get the sin values

OpenStudy (vayacondios1):

2.1?

OpenStudy (peachpi):

I don't know what you're doing. You have to put some work up for me to see. Please start by replacing the letters with the numbers in the equation and we can go from there.

OpenStudy (peachpi):

no. the equation is (sin B)/b = (sin C)/c when you plug in the numbers you should have (sin 45)/3 = (sin 60)/c solve that for c

OpenStudy (vayacondios1):

3.67?

OpenStudy (peachpi):

yes

OpenStudy (vayacondios1):

thank you

OpenStudy (peachpi):

you're welcome